# Optimal Treatment Allocation Strategies for A/B Testing in Partially Observable Time Series Experiments

### v2

@@ Abstract
Time series experiments, in which experimental units receive a sequence of treatments over time, are frequently employed in many technological companies to evaluate the performance of a newly developed policy, product, or treatment relative to a baseline control. Many existing A/B testing solutions assume a fully observable experimental environment that satisfies the Markov condition, which often does not hold in practice.
This paper studies the optimal design for A/B testing in partially observable environments. We introduce a controlled (vector) autoregressive moving average model to capture partial observability. We introduce a small signal asymptotic framework to simplify the analysis of asymptotic mean squared errors of average treatment effect estimators under various designs. We develop two algorithms to estimate the optimal design: one utilizing constrained optimization and the other employing reinforcement learning. We demonstrate the superior performance of our designs using a dispatch simulator and two real datasets from a ride-sharing company. A Python implementation of our proposal is available at https://github.com/datake/ARMADesign.
Keywords: ARMA Model; A/B Testing; Experimental Design; Partially Observable Markov Decision Processes; Policy Evaluation; Reinforcement Learning.
@@ 1 Introduction
Background. A growing number of companies, particularly multi-sided platforms like Airbnb, DoorDash, Uber, and retail marketplaces such as Amazon and Zara are increasingly harnessing data-driven approaches to evaluate and refine their policies and products. In particular, A/B testing, which conducts online experiments to compare a standard control policy “A” to an alternate version “B”, plays a crucial role in informing business decisions within these companies. This method has proven invaluable for their growth and development (Koning et al.,, 2022). For instance, ride-sharing platforms, including Uber, Lyft, and DiDi Chuxing, constantly develop new order dispatching, driver repositioning, and pricing policies. They assess their improvements through A/B testing (Qin et al.,, 2022). Accurate A/B testing enables decision-makers to choose better policies that meet more ride requests, enhance passenger satisfaction, increase driver income, and thus benefit the entire transportation ecosystem (Xu et al.,, 2018; Tang et al.,, 2019).
Challenges. A/B testing in online experiments usually spans a certain duration. This approach presents four major challenges:
@@
    Small sample size. Online experiments are often constrained to a short duration, typically several weeks (Bojinov et al.,, 2023). This limited timeframe leads to large variances in estimating the difference in expected outcomes between the new and standard policies, referred to as the average treatment effect (ATE).
@@
    Carryover effects. Carryover effects are common in time series experiments, where the treatment assigned at a given time can influence future outcomes (Bojinov and Shephard,, 2019; Lewis and Syrgkanis,, 2021; Shi et al.,, 2023; Xiong et al.,, 2023; Chen et al.,, 2024). These effects are typical in ride-sharing companies where past policies can alter the distribution of drivers in the city, which in turn affects future outcomes; refer to Figure 1 for detailed illustrations. Such phenomena often lead to violations of the stable unit treatment value assumption (SUTVA, see Imbens and Rubin,, 2015, Section 1.6), rendering many existing A/B testing solutions (see, e.g., Johari et al.,, 2017; Azevedo et al.,, 2020; Larsen et al.,, 2024; Quin et al.,, 2024; Waudby-Smith et al.,, 2024) and program evaluation methods (see, e.g., Imai and Ratkovic,, 2013; Belloni et al.,, 2017; Chernozhukov et al.,, 2017; Syrgkanis et al.,, 2019; Armstrong and Kolesár,, 2021; Athey et al.,, 2021; Viviano and Bradic,, 2023) ineffective.
@@
    Partial observability. Partial observability frequently occurs in time series experiments. Assuming the underlying time series follows a Markov chain or Markov decision process (MDP, Puterman,, 2014), full observability requires its state to be completely recorded. In contrast, partial observability means only part of the state is observable, leading to the violation of the Markov property (Krishnamurthy,, 2016). It is often the rule rather than the exception in real applications, where recording all relevant features to ensure the ”memoryless” property proves impractical. To elaborate, consider our motivating ride-sharing example. The left panels of Figure 2 visualize two sequences of driver income from two cities, both exhibiting strong daily patterns. The middle and right panels display the auto-correlation function (ACF) and partial ACF (PACF) of the residuals of these income sequences after periodic filtering and regression on other relevant market features, demonstrating the non-Markovian nature of the data.
Figure 2: Visualizations of two sequences of driver income collected from a ride-sharing platform in two cities. Each row plots the data from one of the cities. Left panels: The trend in all drivers’ total income over time. Middle panels: The ACF of the residuals of these income sequences (after the periodic filtering and regression on other relevant market features). Right panels: The PACF of the residuals of these income sequences.
Contributions. Our primary objective is to develop a statistical framework for A/B testing that addresses the above challenges. To tackle the first two challenges, we adopt a design-based approach, focusing on carefully designing the experiment to optimize the data generation process from the online experiments, to minimize the mean squared error (MSE) of the resulting average treatment effect (ATE) estimator. To address the last two challenges, we introduce a controlled (vector) autoregressive moving average ((V)ARMA) model for fitting experimental data. The proposed model is a variant of classical (V)ARMA models (Brockwell and Davis,, 2002, Chapters 3) and represents a rich sub-class of partially observable MDP models (POMDP, see, e.g., Monahan,, 1982). It employs the autoregressive component to accommodate carryover effects and incorporates the moving average error structure to allow for partial observability. Our contributions are listed as follows:
@@
    We introduce the controlled (V)ARMA model for A/B testing to effectively address the challenges of carryover effects and partial observability commonly encountered in time series experiments.
@@
  3. 3.
    We derive two efficiency indicators, which are functions of the model parameters, to compare the statistical efficiencies of three frequently employed designs in estimating the ATE: the alternating-day design, the uniform random design, and the alternating-time design.
  4. 4.
    We propose two innovative algorithms to learn the optimal treatment assignment strategy that minimizes the asymptotic MSE of the resulting ATE estimator: one based on constrained optimization and the other via reinforcement learning (RL).
Our proposal integrates cutting-edge machine learning algorithms, such as RL, with asymptotic theories derived from classical time series models in econometrics, to offer guidance for policy deployment in real-world applications.
Outline. We discuss the related literature in Section 2. In Section 3, we introduce the controlled ARMA model, elaborate its connection with POMDPs, and develop the associated estimating procedure for ATE. We further derive the asymptotic MSEs of different designs under the small signal assumption and propose two efficiency indicators to assess their effectiveness and compare their performance. In Section 4, we present the proposed algorithms to estimate the optimal design. In Section 5, we demonstrate the efficacy of the proposed efficiency indicators and designs through simulations based on a dispatch simulator and two real datasets from a ride-sharing platform. Finally, we conclude our paper in Section 6.
@@ 2 Related Literature
Our proposal intersects with a wide range of research fields, including econometrics, statistics, management science, operational research, and machine learning. It particularly engages with three main research branches: experimental designs, POMDPs, and ARMA and state space models.
@@
Experimental Designs. The design of experiments, also known as experimental design, is a classical problem in statistics, driven by diverse applications in biology, psychology, agriculture, and engineering (Fisher et al.,, 1966). Our proposal specifically relates to papers that focus on identifying treatment allocation strategies tailored for clinical trials (see, e.g., Robbins,, 1952; Begg and Iglewicz,, 1980; Wald,, 2004; Atkinson et al.,, 2007; Jones and Goos,, 2009; Rosenblum et al.,, 2020; Athey and Wager,, 2021). These studies typically focus on non-dynamic settings – often referred to as contextual bandit settings in the machine learning literature – where observations are assumed to be independent, excluding any carryover effects. In contrast, our research on time series experiments accommodates long-term carryover effects and addresses the more complex challenge of temporal dynamics. While traditional crossover designs (Laird et al.,, 1992; Jones and Kenward,, 2003) can deal with long-lasting carryover effects, they often require extended washout periods, making them less practical for modern A/B testing with short durations.
@@
More recently, there has been growing literature that explores experimental designs for A/B testing in technological companies. Our work diverges from them in several key aspects: (i) Many papers do not study time series experiments (Bajari et al.,, 2021; Leung,, 2022; Wan et al.,, 2022; Wang et al., 2023b, ; Basse et al.,, 2024), which is the focus of this paper. (ii) Some existing works adopt an RL framework to model the experimental data, either implicitly (Hu and Wager,, 2022) or explicitly (Li et al.,, 2023; Wen et al.,, 2024), where the data follows a fully observable MDP. In contrast, our framework is more general, accommodating partial observability, which is more typical in real applications. (iii) Several recent studies have shifted focus to switchback designs, where policies alternate at specified intervals under various optimality conditions (Hu and Wager,, 2022; Bojinov et al.,, 2023; Xiong et al.,, 2023; Wen et al.,, 2024). In contrast, our approach considers a broader class of designs that allow each treatment assignment to be influenced by the entire treatment history (see Section 4).
@@
Finally, experimental design is also referred to as the behavior policy search problem in the RL literature, in which Mukherjee et al., (2022); Hanna et al., (2017, 2019) explored the optimal behavior policy by minimizing the MSE of the resulting policy value estimator in MDPs. In contrast, our work focuses on ATE – the difference between two policy value estimators — and allows partial observability, offering a more realistic scenario in practice.
@@
POMDPs. Partial observability often arises in real applications, including autonomous driving (Levinson et al.,, 2011), resource allocation (Bower and Gilbert,, 2005), recommendation (Li et al.,, 2010), and medical management systems (Hauskrecht and Fraser,, 2000). POMDP (Littman and Sutton,, 2001) is the most commonly used model to characterize the partial observability of a stochastic dynamics system. Learning the optimal policy in general POMDPs requires the agent to infer the latent belief state (Krishnamurthy,, 2016), which is both statistically and computationally intractable in general (Papadimitriou and Tsitsiklis,, 1987; Mundhenk et al.,, 2000; Mossel and Roch,, 2005; Vlassis et al.,, 2012). Despite these challenges, it is possible to focus on a sub-class of POMDPs to make the estimation tractable (Kwon et al.,, 2021; Liu et al.,, 2022; Wang et al., 2023a, ). Our proposal follows this principle by introducing a controlled (V)ARMA model under a weak signal condition to streamline estimation. Different from existing works that proposed partial history importance weighting (Hu and Wager,, 2023) or value function-based methods (Uehara et al.,, 2023) to construct policy value estimators, we focus on the experimental design, aiming to optimize the data collection process to enhance policy evaluation.
@@
ARMA and State Space Models. The ARMA model, a cornerstone in time series analysis, has been widely employed in various domains, particularly in econometrics (Hendry,, 1995; Fan and Yao,, 2003; Mikusheva,, 2007, 2012; Box et al.,, 2015; Hamilton,, 2020). Additionally, it is closely related to state space models, which plays a vital role in analyzing continuous dynamic systems (Harvey,, 1990; Durbin and Koopman,, 2012; Aoki,, 2013; Kim and Nelson,, 2017; Komunjer and Zhu,, 2020). The ARMA and state space models are also related to POMDPs, which can be seen as controlled state space models with an added dimension of the action or treatment space, allowing state transitions to be influenced by treatments (Krishnamurthy,, 2016); see Section 3.2 for detailed discussions about their connections.
@@ 3 The Controlled ARMA Model and Its Applications in A/B Testing
This section presents the proposed controlled (V)ARMA model and demonstrates its usefulness in estimating the ATE and comparing distinct treatment allocation strategies. We first describe the data collected from time series experiments, define the ATE for A/B testing, and introduce three commonly used designs in Section 3.1. We further introduce the proposed controlled ARMA model, discuss its connections to POMDPs, and present the estimation procedure for ATE in Section 3.2. Next, we propose the small signal asymptotic framework, establish the asymptotic MSE of the estimated ATE, and then derive two efficiency indicators to compare the estimation efficiency under the three designs in Section 3.3. Finally, we generalize these results to accommodate multivariate observations based on the proposed controlled VARMA model in Section 3.4.
@@ 3.1 Data, ATE, and Designs
  • The subsequent elements of $\mathbf{Y}_{t}$ denote additional relevant market features except $Y_{t,1}$, which can encompass the drivers’ online time and the number of call orders at the $t$-th interval on the online platform in the context of ride-sharing. These features represent the supply and demand of the ride-sharing platform and can significantly influence the outcome (Zhou et al.,, 2021).
@@
provided the limit exists. Here, $\mathbb{E}_{1}$ and $\mathbb{E}_{-1}$ denote expectations under which the treatment $U_{t}$ is consistently set to $1$ and $-1$ at every time $t$, respectively. This objective is a central focus in A/B testing for time series experiments (see, e.g., Hu and Wager,, 2022; Bojinov et al.,, 2023; Li et al.,, 2023; Wen et al.,, 2024). Both terms on the right-hand-side (RHS) of (3.1) should be understood as potential outcomes (Imbens and Rubin,, 2015), representing the average outcome that would have been observed if either the new treatment or the control had been assigned at all times. Nonetheless, as we focus on experimental design, it eliminates concerns about unmeasured confounders. To simplify the presentation, we choose not to use potential outcome notations. Interested readers may refer to Ertefaie, (2014), Bojinov and Shephard, (2019), Luckett et al., (2020), Shi et al., (2023), and Viviano and Bradic, (2023) for detailed discussions on potential outcomes in dynamic settings.
@@
Design. In our context, each design corresponds to a sequence of treatment allocation strategies $\pi=\{\pi_{t}\}_{t\geq 1}$ where each $\pi_{t}$ specifies the conditional distribution of $U_{t}$ given the past data history up to time $t-1$, denoted by $\mathbf{H}_{t-1}=\{\mathbf{Y}_{1},U_{1},\ldots,\mathbf{Y}_{t-1},U_{t-1}\}$. Informally speaking, each design determines the probabilities of applying the treatment and control at each time, given the past history. Our focus is on observation-agnostic designs, where each $\pi_{t}$ depends on $\mathbf{H}_{t-1}$ only through $\{U_{1},U_{2},\ldots,U_{t-1}\}$, independent of past observations. In this section, we specifically consider the following three designs within this category:
  1. 1.
@@
    Alternating-time (AT) design: This design alternates between treatment and control at adjacent time intervals and is frequently employed in many ride-sharing companies, such as Lyft and DiDi Chuxing, to compare different order-dispatching policies (Chamandy,, 2016; Luo et al.,, 2024). To implement the AT design, the initial treatment $U_{1}$ is randomly generated with equal probabilities: $\pi_{1}(1)=\pi_{1}(-1)=0.5$. For subsequent times, we set $\pi_{t}(-U_{t-1}|\mathbf{H}_{t-1})=1$ and $\pi_{t}(U_{t-1}|\mathbf{H}_{t-1})=0$, ensuring that $U_{t}=-U_{t-1}$ almost surely.
  2. 2.
@@
    Alternating-day (AD) design: This design assigns the same treatment throughout each day and switches to the control or the opposite treatment on the following day. Similar to the AT design, the initial treatment $U_{1}$ in the AD design is also uniformly randomly determined, with $\pi_{1}(1)=\pi_{1}(-1)=0.5$. Let $\tau$ represent the number of time intervals per day. The treatment assignment ensures that $U_{1}=U_{2}=\cdots=U_{\tau}=-U_{\tau+1}=-U_{\tau+2}=\cdots=-U_{2\tau}=U_{2\tau+1}=\cdots$, maintaining consistency within each day and alternating on a daily basis.
  3. 3.
@@
    Uniform random (UR) design: This design independently assigns treatment and control randomly with equal probabilities each time. Specifically, $\pi_{t}$ remains a constant function with a value of 0.5, regardless of $t$ and $\mathbf{H}_{t-1}$. Despite its simplicity, designs of this type have been widely adopted in clinical trials (Almirall et al.,, 2014).

### v4

# JRSSB_supplementary

@@ Abstract
Online experiments are frequently employed in many technological companies to evaluate the performance of a newly developed policy, product, or treatment relative to a baseline control. In many applications, the experimental units receive a sequence of treatments over time. To handle these time-dependent settings, existing A/B testing solutions typically assume a fully observable experimental environment that satisfies the Markov condition. However, this assumption often does not hold in practice.
This paper studies the optimal design for A/B testing in partially observable online experiments. We introduce a controlled (vector) autoregressive moving average model to capture partial observability. We introduce a small signal asymptotic framework to simplify the calculation of asymptotic mean squared errors of average treatment effect estimators under various designs. We develop two algorithms to estimate the optimal design: one utilizing constrained optimization and the other employing reinforcement learning. We demonstrate the superior performance of our designs using two dispatch simulators that realistically mimic the behaviors of drivers and passengers to create virtual environments, along with two real datasets from a ride-sharing company. A Python implementation of our proposal is available at https://github.com/datake/ARMADesign.
Keywords: ARMA Model; A/B Testing; Experimental Design; Partially Observability; Policy Evaluation; Reinforcement Learning.
@@ 1 Introduction
Background. A growing number of companies, particularly multi-sided platforms like Airbnb, DoorDash, Uber, and retail marketplaces such as Amazon and Zara are increasingly harnessing data-driven approaches to evaluate and refine their policies and products. In particular, A/B testing, which conducts online experiments to compare a standard control policy “A” to an alternate version “B”, plays a crucial role in informing business decisions within these companies and has proven invaluable for their growth and development (Koning et al.,, 2022). For instance, ride-sharing platforms, including Uber, Lyft, and DiDi Chuxing, constantly develop new order dispatching, driver repositioning, pricing policies and assess their improvements through A/B testing (Qin et al.,, 2024). Accurate A/B testing enables decision-makers to choose better policies that meet more ride requests, enhance passenger satisfaction, increase driver income, and thus benefit the entire transportation ecosystem (Xu et al.,, 2018).
Challenges. In many applications, the experimental units receive treatments sequentially over time. A/B testing in these experiments poses four major challenges:
@@
    Small sample size. Online experiments are often constrained to a short duration, typically several weeks (Luo et al.,, 2024). This limited timeframe leads to large variances in estimating the difference in expected outcomes between the new and standard policies, referred to as the average treatment effect (ATE).
@@
    Carryover effects. Carryover effects are ubiquitous in online experiments, where the treatment assigned at a given time can influence future outcomes (Bojinov and Shephard,, 2019; Han et al.,, 2022; Shi et al., 2023b, ; Xiong et al.,, 2023; Chen et al.,, 2024). These effects are typical in ride-sharing companies where past policies can alter the distribution of drivers in the city, which in turn affects future outcomes; refer to Figure 1 for detailed illustrations. Such phenomena lead to violations of the stable unit treatment value assumption (SUTVA, see Imbens and Rubin,, 2015, Section 1.6), rendering many existing A/B testing solutions (see, e.g., Johari et al.,, 2017; Azevedo et al.,, 2020; Wang et al.,, 2023; Larsen et al.,, 2024; Quin et al.,, 2024; Waudby-Smith et al.,, 2024) and causal inference methods (see, e.g., Imai and Ratkovic,, 2013; Belloni et al.,, 2017; Chernozhukov et al.,, 2017; Armstrong and Kolesár,, 2021; Athey et al.,, 2021; Viviano and Bradic,, 2023; Ding,, 2024) ineffective.
@@
    Figure 2: Visualizations of two sequences of driver income collected from a ride-sharing platform in two cities. Each row plots the data from one of the cities. Left panels: The trend in all drivers’ total income over time. Middle panels: The ACF of the residuals of these income sequences (after filtering the seasonal effects within each day and regressing on other relevant market features). Right panels: The PACF of the residuals of these income sequences.
@@
    Partial observability. Partial observability frequently occurs in online experiments. Assuming the underlying time series follows a Markov chain or Markov decision process (MDP, Puterman,, 2014), full observability requires its state to be completely recorded. In contrast, partial observability means only part of the state is observable, leading to the violation of the Markov property (Krishnamurthy,, 2016). It is often the rule rather than the exception in real applications, where recording all relevant features to ensure the “memoryless” property proves impractical. To elaborate, consider our motivating ride-sharing example. The left panels of Figure 2 visualize two sequences of driver income from two cities, both exhibiting strong daily patterns. The middle and right panels display the auto-correlation function (ACF) and partial ACF (PACF) of the residuals of these income sequences after filtering the seasonal effects within each day and regressing on other relevant market features. Notably, the PACF exhibits significant higher-order lags, which demonstrates the non-Markovian nature of the data.
Contributions. Our primary objective is to develop a statistical framework for A/B testing that addresses the above challenges. Our contributions include the development of optimal designs, efficiency indicators, statistical modeling, estimation methods, and theoretical frameworks. We detail them as follows.
@@
    To tackle the first two challenges, we focus on carefully designing the experiment to optimize the data generation process from the online experiments, so as to minimize the mean squared error (MSE) of the resulting average treatment effect (ATE) estimator. In particular, we propose two innovative algorithms to learn the optimal design: one based on constrained optimization and the other via reinforcement learning (RL). Our empirical studies, which leverage a synthetic dispatch simulator and a city-level real-data-based simulator – both constructed using physical models to realistically simulate driver and passenger behaviors – along with two real datasets from a ride-sharing company, demonstrate that our proposed designs consistently outperform existing state-of-the-art.
@@
    Additionally, we derive two efficiency indicators to compare the statistical efficiencies of three frequently employed designs in estimating the ATE: the alternating-day design, the uniform random design, and the alternating-time design.
  3. 3.
    To address the last two challenges, we introduce a controlled (vector) autoregressive moving average ((V)ARMA) model for fitting experimental data. The proposed model is a variant of classical (V)ARMA models (Brockwell and Davis,, 2002, Chapters 3) and represents a rich sub-class of partially observable MDP models (POMDP, see, e.g., Monahan,, 1982). It employs the autoregressive component to accommodate carryover effects and incorporates the moving average error structure to allow for partial observability.
  4. 4.
@@
To summarize, our proposal integrates cutting-edge machine learning algorithms, such as RL, with asymptotic theories derived from classical time series models in econometrics to offer guidance for policy deployment in real-world applications.
Outline. We discuss the related literature in Section 2. In Section 3, we introduce the controlled ARMA model, elaborate its connection with POMDPs, and develop the associated estimating procedure for ATE. We further derive the asymptotic MSEs of different designs under the small signal assumption and propose two efficiency indicators to assess their effectiveness. In Section 4, we present the proposed algorithms to estimate the optimal design. In Section 5, we demonstrate the efficacy of the proposed designs and efficiency indicators through two dispatch simulators and two real datasets from a ride-sharing platform. Finally, we conclude our paper in Section 6.
@@ 2 Related Literature
Our proposal intersects with a wide range of research fields, including econometrics, statistics, management science, operational research, and machine learning. It particularly engages with three main research branches: experimental designs, POMDPs, ARMA and state space models.
@@
Experimental Designs. The design of experiments, also known as experimental design, is a classical problem in statistics, driven by diverse applications in biology, psychology, agriculture, and engineering (Fisher et al.,, 1966). Within the statistics literature, our proposal specifically relates to works that focus on identifying treatment allocation strategies tailored for clinical trials (see, e.g., Robbins,, 1952; Pocock and Simon,, 1975; Begg and Iglewicz,, 1980; Atkinson et al.,, 2007; Jones and Goos,, 2009; Rosenblum et al.,, 2020; Liu and Hu,, 2022; Ma et al.,, 2024). These studies typically focus on non-dynamic settings (referred to as contextual bandit settings in the machine learning literature) where observations are assumed to be independent, excluding any carryover effects. In contrast, our research accommodates carryover effects and addresses the more complex challenge of temporal dynamics. While traditional crossover designs (Laird et al.,, 1992; Jones and Kenward,, 2003) can deal with long-lasting carryover effects, they often require extended washout periods, making them less practical for modern A/B testing with short durations.
@@
More recently, there has been a growing body of literature in management science, econometrics, and machine learning that explores experimental designs for A/B testing in technological companies. Our work differs from them in several aspects: (i) Many papers consider settings without carryover effects over time (Bajari et al.,, 2021; Wan et al.,, 2022; Viviano et al.,, 2023; Wang et al.,, 2023; Basse et al.,, 2024). (ii) Some existing works adopt an RL framework to model the experimental data (Glynn et al.,, 2020; Li et al.,, 2023; Wen et al.,, 2024), where the data follows a fully observable MDP. In contrast, our framework is more general and accommodates partial observability, which is a more typical scenario in real applications. (iii) Several recent studies focus on switchback designs where policies alternate at specified intervals under various optimality conditions (Hu and Wager,, 2022; Bojinov et al.,, 2023; Xiong et al.,, 2023; Wen et al.,, 2024). In contrast, our approach considers a broader class of designs that allow each treatment assignment to be influenced by the entire treatment history (see Section 4).
@@
In the RL literature, the design of the experiment is also referred to as the behavior policy search problem, in which Mukherjee et al., (2022); Hanna et al., (2017) explored the optimal behavior policy by minimizing the MSE of the policy value estimator in MDPs. Meanwhile, Agarwal et al., (2022, Section 3.3) employed the D-optimal design for policy learning in MDPs. In contrast to these works, we focus on the evaluation of ATE – the difference between two policy value estimators — and allow partial observability, offering a more realistic scenario in practice.
Finally, recent works have approached the design problem from an optimization perspective (Zhao,, 2024). In particular, works in the machine learning literature have proposed the use of deep learning or RL to numerically compute a Bayesian version of the optimal design (Foster et al.,, 2021; Blau et al.,, 2022). In contrast to these methods, we employ a frequentist approach and focus specifically on the evaluation of ATE.
@@
POMDPs. Partial observability often arises in real applications, including autonomous driving (Levinson et al.,, 2011), resource allocation (Bower and Gilbert,, 2005), recommendation (Li et al.,, 2010), and medical management systems (Hauskrecht and Fraser,, 2000). POMDP is the most commonly used model to characterize the partial observability of a stochastic dynamics system. Learning the optimal policy in general POMDPs requires the agent to infer the latent belief state (Krishnamurthy,, 2016), which is both statistically and computationally intractable in general (Papadimitriou and Tsitsiklis,, 1987; Vlassis et al.,, 2012). Despite these challenges, it is possible to focus on a sub-class of POMDPs to make the estimation tractable (Kwon et al.,, 2021; Liu et al.,, 2022). Our proposal follows this principle by introducing a controlled (V)ARMA model under a weak signal condition to streamline estimation and design. Different from existing works that proposed partial history importance weighting (Hu and Wager,, 2023) or value-function-based methods (Uehara et al.,, 2023) to construct policy value estimators, we focus on the experimental design, aiming to optimize the data collection process to enhance policy evaluation.
@@
ARMA and State Space Models. The ARMA model, a cornerstone in time series analysis, has been widely employed in various domains, particularly in econometrics (Brockwell,, 1991; Hendry,, 1995; Fan and Yao,, 2003; Box et al.,, 2015; Hamilton,, 2020). Additionally, it is closely related to state space models, which plays a vital role in analyzing continuous dynamic systems (Harvey,, 1990; Durbin and Koopman,, 2012; Aoki,, 2013; Kim and Nelson,, 2017; Komunjer and Zhu,, 2020). The ARMA and state space models are also related to POMDPs, which can be seen as controlled state space models with an added dimension of the action or treatment space, allowing state transitions to be influenced by treatments (Krishnamurthy,, 2016); see Section 3.2 for detailed discussions about their connections.
In the causal inference literature, Menchetti et al., (2021) proposed causal versions of ARIMA models. More recently, Liang and Recht, (2023) proposed linear state space models for causal inference. Despite the similarity in the models, these works differ from ours primarily in their focus: they focused on estimating and inferring causal effects, whereas we concentrate on the experimental design to “optimize” the estimated causal effect. Consequently, these works did not utilize the small signal framework we propose to derive closed-form solutions for the asymptotic MSEs under different designs, which we use as criteria to optimize the design. Nor did they explore RL approaches to finding the optimal design. This difference in objectives also influences the choice of models. For instance, the causal ARIMA model is designed for settings with a single persistent treatment over time (Menchetti et al.,, 2021, Assumption 1), making it unsuitable for studying general designs.
@@ 3 The Controlled ARMA Model and Its Applications in A/B Testing
This section presents the proposed controlled (V)ARMA model and demonstrates its usefulness in estimating the ATE and comparing different treatment allocation strategies. We first describe the data collected from time series experiments, define the ATE for A/B testing, and introduce three commonly used designs in Section 3.1. We further introduce the proposed controlled ARMA model, discuss its connections to POMDPs, and present the estimation procedure for ATE in Section 3.2. Next, we propose the small signal asymptotic framework, establish the asymptotic MSE of the estimated ATE, and then derive two efficiency indicators to compare the estimation efficiency under the three designs in Section 3.3. Finally, we generalize these results to accommodate multivariate observations and exogenous variables based on the proposed controlled VARMA model in Section 3.4.
@@ 3.1 Data, ATE, and Designs
  • The subsequent elements of $\mathbf{Y}_{t}$ denote additional relevant market features, which can contain the drivers’ online time and the number of call orders at the $t$-th interval on the online platform in the context of ride-sharing. These features represent the supply and demand of the ride-sharing platform and can significantly influence the outcome (Zhou et al.,, 2021). Our experiments suggest that jointly modeling both the outcome and market features can substantially improve the estimation of the ATE, achieving a reduction in the MSE by 10 to 100 times compared to approaches that model only the outcome, such as those in Menchetti et al., (2021) and Liang and Recht, (2023).
@@
provided the limit exists. Here, $\mathbb{E}_{1}$ and $\mathbb{E}_{-1}$ denote expectations under which the treatment $U_{t}$ is consistently set to $1$ and $-1$ at every time $t$, respectively. This objective is a central focus in A/B testing with carryover effects (see, e.g., Hu and Wager,, 2022; Li et al.,, 2023; Xiong et al.,, 2023; Wen et al.,, 2024). Both terms on the right-hand-side (RHS) of (3.1) should be understood as potential outcomes (Imbens and Rubin,, 2015), representing the average outcome that would have been observed if either the new treatment or the control had been assigned at all times. Nonetheless, as we focus on experimental design, it eliminates concerns about unmeasured confounders. To simplify the presentation, we choose not to use potential outcome notations. Interested readers may refer to Ertefaie, (2014), Luckett et al., (2020) and Viviano and Bradic, (2023) for detailed discussions on potential outcomes in dynamic settings.
@@
Design. In our context, each design corresponds to a sequence of treatment allocation strategies $\pi=\{\pi_{t}\}_{t\geq 1}$ where each $\pi_{t}$ specifies the conditional distribution of $U_{t}$ given the past data history up to time $t-1$, denoted by $\mathbf{H}_{t-1}=\{\mathbf{Y}_{1},U_{1},\ldots,\mathbf{Y}_{t-1},U_{t-1}\}$. Informally speaking, each design determines the probabilities of applying the treatment and control at each time, given the past history. Our focus is on observation-agnostic designs, where each $\pi_{t}$ depends on $\mathbf{H}_{t-1}$ only through $\{U_{1},U_{2},\ldots,U_{t-1}\}$, independent of past observations. This class covers the following three special examples:
@@
Example I: Alternating-time (AT) design. This design alternates between treatment and control at adjacent time intervals and is frequently employed in many ride-sharing companies, such as Lyft and DiDi Chuxing, to compare different order dispatching policies (Chamandy,, 2016; Luo et al.,, 2024). To implement the AT design, the initial treatment $U_{1}$ is randomly generated with equal probabilities: $\pi_{1}(1)=\pi_{1}(-1)=0.5$. For subsequent times, we set $\pi_{t}(-U_{t-1}|\mathbf{H}_{t-1})=1$ and $\pi_{t}(U_{t-1}|\mathbf{H}_{t-1})=0$ so that $U_{t}=-U_{t-1}$ almost surely.
@@
Example II: Alternating-day (AD) design. This design assigns the same treatment throughout each day and switches to the opposite treatment on the following day. Similar to the AT design, the initial treatment $U_{1}$ in the AD design is also uniformly randomly determined. Let $\tau$ represent the number of time intervals per day. The treatment assignment ensures that $U_{1}=U_{2}=\cdots=U_{\tau}=-U_{\tau+1}=-U_{\tau+2}=\cdots=-U_{2\tau}=U_{2\tau+1}=\cdots$, maintaining consistency within each day and alternating on a daily basis.
@@
Example III: Uniform random (UR) design. This design independently assigns treatment and control randomly with equal probabilities each time. Specifically, $\pi_{t}$ remains a constant function with a value of 0.5, regardless of $t$ and $\mathbf{H}_{t-1}$. Despite its simplicity, designs of this type have been widely adopted in clinical trials.